OpenStudy (fools101):

Rules & Application of Exponents (8).^2/3 Rules & Application of Exponents (8).^2/3 @Mathematics

OpenStudy (fools101):

please help!

OpenStudy (anonymous):



OpenStudy (anonymous):

XD

OpenStudy (anonymous):

8 ^(2/3) = (8^2)^(1/3) =(64)^(1/3)=(4*4*4)(1/3) =4

OpenStudy (anonymous):

= (cuberoot 8 ) ^2 = 2^2 = 4

OpenStudy (anonymous):

When raising to a fractional exponent, the numerator is the exponent on the number or variable. The denominator is the root. So it would be the third/cubed root of 8^2. Or the cubed root of 64. You can do it either way. You can first square 8, or take the cubed root of it. In this case it's easier to first take the cubed root, then squaring the result.

OpenStudy (fools101):

kk but would it come out to be 1/4?

OpenStudy (anonymous):

^(1/3) = cube root

OpenStudy (anonymous):

no it is 2^2 = 4

OpenStudy (fools101):

oohh kk just making sure thank u Jimmyrep : )